# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

### v1

@@ Abstract
Existing point cloud representation learning tend to learning the geometric distribution of objects through data-driven approaches, emphasizing structural features while overlooking the relationship between the local information and the whole structure. Local features reflect the fine-grained variations of an object, while the whole structure is determined by the interaction and combination of these local features, collectively defining the object’s shape. In real-world, objects undergo elastic deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric properties. Inspired by this, we propose a physics-driven self-supervised learning method for point cloud representation, which captures the relationship between parts and the whole by constructing a local-whole force propagation mechanism. Specifically, we employ a dual-task encoder-decoder framework, integrating the geometric modeling capability of implicit fields with physics-driven elastic deformation. The encoder extracts features from the point cloud and its tetrahedral mesh representation, capturing both geometric and physical properties. These features are then fed into two decoders: one learns the whole geometric shape of the point cloud through an implicit field, while the other predicts local deformations using two specifically designed physics information loss functions, modeling the deformation relationship between local and whole shapes. Experimental results show that our method outperforms existing approaches in object classification, few-shot learning, and segmentation, demonstrating its effectiveness.
1 Introduction
@@
Point clouds, as a common data structure, provide a compact representation of 3D objects, capable of precisely describing the geometric shape and spatial distribution of objects. Due to their rich geometric information and broad range of applications, point clouds play a crucial role in 3D vision tasks, such as 3D shape analysis [52], object detection [38], and segmentation [15, 8]. The success of these tasks depends on the accurate modeling of point cloud features, while the irregularity and sparsity of point clouds make learning their effective representations a challenging but essential task.
Figure 1: Deformation visualization. When different magnitudes of forces $({f_{1}}<{f_{2}})$ are applied from top to bottom on the hat, it undergoes varying degrees of deformation.
@@
Existing methods primarily rely on data-driven approaches to capture geometric distributions and structural features of point clouds. In supervised point cloud representation learning, typical network architectures, such as PointNet [30], use shared MLPs to process unordered point sets and directly learn the feature representations of each point in the point cloud to capture the whole geometric distribution. PointNet++ [31] builds upon this by introducing a hierarchical structure to capture local features, allowing for better capture of structural information in local regions. DGCNN [45] utilizes dynamic graph convolutional networks to capture local geometric structural features. In unsupervised point cloud representation learning, generative methods [56, 55, 23, 13] learn latent representations of the data by reconstructing the geometric structure of the point cloud using generative models such as autoencoder (AE) [24], generative adversarial networks (GAN) [7], or autoregressive (AR) [39, 23]. Discriminative methods [49, 16, 47, 22] typically construct tasks by comparing feature differences between different augmented samples or categories, focusing more on learning the whole distribution patterns [14, 34] of point clouds in the feature space. Despite the significant progress made by these methods in point cloud representation learning, they often overlook the interaction between local information and whole structure, failing to explicitly model their interrelationship.
@@
Specifically, existing methods often expand local information to the whole level through multi-scale operations (e.g., the hierarchical structure of PointNet++) or max-pooling operations (e.g., global feature aggregation in PointNet). This process of aggregating information solely in the feature space does not explicitly model and capture the interrelationship between local and whole structures. For example, DGCNN captures local geometric structures using dynamic graph convolutional networks, but its layer-by-layer expansion of local information remains a passive, lacking explicit modeling of the relationship between part and whole structural features.
@@
In the real world, objects undergo elastic deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric properties. As shown in Figure 1, objects change to different degrees when forces of different magnitudes are applied to them. Based on this, we propose a physics-driven self-supervised learning method for point cloud representation, which captures the deformation effects between local and whole shapes by introducing the force propagation mechanism from elasticity theory. Specifically, we design a dual-task encoder-decoder framework. The encoder is responsible for learning the geometric information and physical properties of the point cloud from both the point cloud data and its corresponding tetrahedral mesh representation. The learned features are then input into two decoders: one decoder models the whole geometric shape of the point cloud using implicit fields, while the other decoder uses data-fidelity loss function and physics-informed loss function to predict local deformations, simulating the force propagation relationship between local and global structures. The main contributions of our method are as follows:
@@
$\bullet$ We propose a novel self-supervised learning method that explicitly captures the relationship between fine-grained features and whole structure in point clouds by combining physical laws with geometric modeling. This is also the first method to apply elastic mechanics to self-supervised learning.
@@
$\bullet$ We demonstrate the effectiveness of our method through extensive evaluations on multiple downstream tasks, including object classification, few-shot learning, and segmentation. Additionally, the incorporation of physical information enhances the understanding of 3D point clouds.
2 Related Works
Point Cloud Representation Learning: In recent years, the advancement of deep learning techniques has enabled direct utilization of point cloud data for learning intrinsic feature representations of shapes. Early methods, such as PointNet [30], used shared MLPs and max pooling operations to learn global features of point clouds. Subsequently, PointNet++ [31] introduced a hierarchical structure and local feature learning mechanisms to address the issue of PointNet’s insensitivity to local features in point clouds. DGCNN [45] further improves local feature extraction by dynamically constructing graphs based on point neighborhoods. Similarly, recent studies [32, 33, 19] have focused on learning local feature information from point clouds, aiming to extract as many semantic features as possible. However, the methods often treat point clouds as static geometric distributions, overlooking the explicit relationship between local and whole structures. This limitation, especially in the case of object deformation, makes it difficult to capture the impact of local changes on the whole structure.
@@ 2 Related Works
Unsupervised Point Cloud Representation Learning: Nowadays, numerous unsupervised learning methods have emerged. Generative methods learn features through self-reconstruction [44, 56, 20, 50, 13, 41]. For example, HC-VAE [20] introduces a hierarchical consistency mechanism to progressively generate high-quality point clouds. FoldingNet [50] reconstructs arbitrary point clouds using a graph-based encoder and a folding-based decoder. Recently, recovering missing parts from masked inputs [55, 23, 56, 53] has also become a prominent research focus. MaskSurf [55] designs a transformer-based encoder-decoder network to estimate the position of the masking point and its corresponding normal vector. Point-DAE [56] reconstructs various corrupted point clouds by decomposing the reconstruction of a complete point cloud into detailed local patches and a coarse global shape, thereby alleviating the issue of positional leakage in the reconstruction process. However, generative methods tend to learn the latent representations of point clouds through whole reconstruction tasks, overlooking local features and their impact on the whole structure.
Discriminative methods guide the model to learn meaningful feature representations by designing tasks that focus on the feature differences between samples. Some methods employ pretext tasks to learn point cloud representations [36, 37, 49, 46, 26, 4]. Jigsaw3D [36] uses 3D puzzles as a self-supervised learning task, allowing the model to learn how to correctly assemble object parts. Poursaeed et al. [29] define the estimation of point cloud rotation angles as a pretext task. Recently, contrastive learning methods [16, 21, 49] have been widely used in 3D self-supervised learning, achieving state-of-the-art performance. PointContrast [49] performs point-level invariant mapping on two different directional views of the same point cloud. It is the first unified framework to study the contrastive paradigm for 3D representation learning. STRL [16] is an extension of the method [11] on 3D point clouds to learn rich representational information by updating between online and target networks. PointSmile [22] guides network learning by maximizing the mutual information between two point clouds. However, discriminative methods typically rely on a large number of sample pairs and carefullydesigned learning strategies, making it challenging to effectively learn the interactions between local deformations and global shapes.
In addition, some researchers have focused on incorporating cross-model [1, 47, 54, 18, 28] methods into discriminative methods. CrossPoint [1] integrates both 2D and 3D modal information, generating more robust point cloud features. Building upon CrossPoint, CrossNet [47] further divides the images into RGB and grayscale for comparative analysis. TCMSS [12] parameterizes point cloud and image data as Dirichlet distributions and then aligns the data from different modalities.
3 Methods
3.1 Overview
@@
The framework of our proposed method is illustrated in Figure 2. It is primarily composed of two modules: The first module is the implicit feature learning module, where the point cloud is input into an autoencoder to learn its implicit features. The implicit field represents the point cloud as a continuous surface, and by leveraging the distance function from a query point to the nearest surface, it can uniformly handle objects of arbitrary shapes, helping the network better learn the whole shape features of the object. The second module is the physical information awareness module, which captures the mechanical characteristics of local deformations in the point cloud to model the relationship between local features and the whole geometric shape. Our autoencoder-based dual-task learning framework not only provides a more comprehensive representation of geometric structures but also incorporates physical constraints, enabling more efficient learning of point cloud representations with real-world significance.
3.2 Implicit Feature Learning
@@
For a point cloud dataset $\mathcal{D}=\left\{{{P}_{i}}\right\}_{i=1}^{\left|\mathcal{D}\right|}$, we use encoder ${{\Phi}_{\theta}}$ and decoder ${{\Psi}_{\eta}}$ to learn the implicit features of the point cloud ${P_{i}}=\left\{{p_{i}^{j}\in{\mathbb{R}^{3}}|j=1,2,...,n}\right\}$. The encoder ${{\Phi}_{\theta}}$ takes the ${P_{i}}$ as input and maps it to an $m$-dimensional latent vecoter ${z_{\eta}}$, represented as ${{\Phi}_{\theta}}:{{\mathbb{R}}^{n\times 3}}\to{{\mathbb{R}}^{m}},m\ll n$. Where we use the network backbones of PointNet [30] and DGCNN [45] as ${{\Phi}_{\theta}}$ to learn the point cloud features, respectively. Subsequently, we concatenate latent vector ${z_{\eta}}$ with an arbitrary query point $q$ in 3D space to obtain feature vector $z_{\eta}^{q}={z_{\eta}}\oplus q$. Then, the decoder ${{\Psi}_{\eta}}$ takes $z_{\eta}^{q}$ as input and output an implicit function $\mathbb{S}={{\Psi}_{\eta}}\left(q,{{\Phi}_{\theta}}\left({{P}_{i}}\right)\right)=\left\{{\delta_{i}}|{\delta_{i}}\left({z_{\eta}^{q}}\right)\in\mathbb{R},\forall q\in{{\mathbb{R}}^{3}}\right\}$. ${\delta_{i}}\left({z_{\eta}^{q}}\right)$ represents the shortest distance from the query point $q$ to the point cloud ${P_{i}}$. The decoder ${{\Psi}_{\eta}}$ consists of 3-layer of MLP and outputs a 1-dimensional feature vector.
@@
${\theta^{*}},{\eta^{*}}=\mathop{\arg\min}\limits_{\theta,\eta}\left({\delta_{i}(z_{\eta}^{q})),{\delta_{gt}}\left(q\right)}\right)$ (1)
@@
${{\cal L}_{imp}}=\frac{1}{{|\mathcal{D}|K}}\sum\limits_{i=1}^{|\mathcal{D}|}{\sum\limits_{q\in Q}^{K}{\left|{\left|{{\delta_{i}}(z_{\eta}^{q})}\right|-{\delta_{gt}(q)}}\right|}}$ (2)
@@
3.3 Physical Information Awareness
For a point cloud ${{P}_{i}}$, we first use 3D Delaunay [27] to convert the ${{P}_{i}}$ into an object composed of multiple tetrahedra. It is an extension of Delaunay triangulation (which generates triangular meshes in 2D) to 3D space, producing a mesh composed of tetrahedral elements. Compared to hexahedral meshes, tetrahedral meshes are able to fit free-form objects efficiently and are well-suited for complex topological structures. The meshing process of the point cloud is illustrated in Figure 3. Subsequently, the tetrahedral mesh data is processed by the mesh processor $\mathrm{T}$ to be mapped into a format suitable for input to encoder ${{\Phi}_{\theta}}$. Where the mesh processor $\mathrm{T}$ consists of 3-layer of convolutional networks, serving as a bridge between the mesh data and encoder ${{\Phi}_{\theta}}$. The encoder ${{\Phi}_{\theta}}$ maps the processed data to the $m$-dimensional latent space to obtain the latent vector ${z_{phys}}$. ${z_{phys}}$ is concatenated with the deformation parameters $d$ as the input $z_{phy}^{d}$ to decoder ${{\Psi}_{phys}}$, enabling ${{\Psi}_{phys}}$ to predict the displacement field $\hat{u}$ of the object. Decoder ${{\Psi}_{phys}}$ is composed of 5-layer of MLP, outputting a feature vector of size $n\times 3$, where $n$ is the number of input points in the point cloud.

### v2

@@ Abstract
Existing point cloud representation learning methods primarily rely on data-driven strategies to extract geometric information from large amounts of scattered data. However, most methods focus solely on the spatial distribution features of point clouds while overlooking the relationship between local information and the whole structure, which limits the accuracy of point cloud representation. Local information reflect the fine-grained variations of an object, while the whole structure is determined by the interaction and combination of these local features, collectively defining the object’s shape. In real-world, objects undergo deformation under external forces, and this deformation gradually affects the whole structure through the propagation of forces from local regions, thereby altering the object’s geometric features. Therefore, the appropriate introduction of physics-driven mechanism can effectively compensate for the limitations of data-driven methods in structural modeling and significantly enhance the generalization and interpretability of point cloud representations in downstream tasks such as understanding and recognition. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method to learn fine-grained features in point clouds and model the structural relationship between local regions and the whole shape. Specifically, we design a dual-task encoder-decoder framework that combines the geometric modeling capability of data-driven implicit fields with physics-driven elastic deformation. Through the integration of physics-based loss functions, the framework is guided to predict localized deformation and explicitly capture the correspondence between local structural changes and whole shape variations. Experimental results show that our method outperforms existing approaches in object classification and segmentation, demonstrating its effectiveness.
Index Terms:
Point Cloud Representation, Physics-driven, Implicit Field, Self-supervised Learning.
I Introduction
@@
Point clouds can accurately describe the geometric shape and spatial distribution of objects. They are widely used in tasks such as 3D shape analysis [1], object detection [2], and segmentation [3, 4]. The success of these tasks depends on the accurate modeling of point cloud features, while the irregularity and sparsity of point clouds make learning their effective representations a challenging but essential task.
Figure 1: (a) Visualization of deformation and force distribution under an external force $f$ for Lamp A and B. Colors represent force magnitude, with red circles indicating regions of significant deformation. (b) Clustering results with and without Physics-Informed (PI) modeling.
@@
Existing methods primarily rely on data-driven approaches to capture geometric distributions of point clouds. In supervised point cloud representation learning, typical network architectures, such as PointNet [5], use shared MLPs to process unordered point sets and directly learn the feature representations of each point in the point cloud to capture the whole geometric distribution. PointNet++ [6] builds upon this by introducing a hierarchical structure to capture local features, allowing for better capture of structural information in local regions. DGCNN [7] utilizes dynamic graph convolutional networks to capture local geometric structural features. In unsupervised point cloud representation learning, generative methods [8, 9, 10, 11] learn latent representations of the data by reconstructing the geometric structure of the point cloud using generative models such as autoencoder (AE) [12], generative adversarial networks (GAN) [13], or autoregressive (AR) [10]. Discriminative methods [14, 15, 16, 17] typically construct tasks by comparing feature differences between different augmented samples or categories, focusing more on learning the whole distribution patterns [18, 19] of point clouds in the feature space. Despite the significant progress made by these methods in point cloud representation learning, they often overlook the interaction between local information and whole structure, failing to explicitly model their interrelationship.
@@
Specifically, existing methods often expand local information to the global level through multi-scale operations (e.g., the hierarchical structure of PointNet++) or max-pooling operations (e.g., global feature aggregation in PointNet). This process of aggregating information solely in the feature space does not explicitly model and capture the interrelationship between local and whole structures. For example, DGCNN captures local geometric structures using dynamic graph convolutional networks, but its layer-by-layer expansion of local information remains a passive, lacking explicit modeling of the relationship between part and whole structural features.
@@
In nature, the shape of objects is often governed by the forces acting upon them. When an object is subjected to external forces, slight deformations first occur in its local regions. These deformations propagate through structural connections, causing global shape changes. This transition from “local stress” to “global response” reveals the inherent coupling relationships among the parts of an object. Meanwhile, recent research [20] has shown that physical modeling not only provides interpretability and inductive capability, but can also be embedded into deep networks as structural constraints to enhance the performance of data-driven methods on downstream detection tasks. As shown in Figure 1, incorporating the physics-driven approach enables more effective discrimination of fine-grained features both intra-class and inter-class. Inspired by this, we incorporate a physics-driven mechanism into the data-driven method, which captures the deformation effects between local and whole shapes by introducing the force propagation mechanism from elasticity theory. Specifically, we design a dual-task encoder-decoder framework. The encoder is responsible for learning the geometric information and physical properties of object from both the point cloud data and its corresponding tetrahedral mesh representation. The learned features are then input into two decoders: one decoder models the whole geometric shape of the point cloud using implicit fields, while the other decoder uses data-fidelity loss function and physics-informed loss function to predict local deformations, simulating the force propagation relationship between local and whole structures. The main contributions of our method are as follows:
@@
$\bullet$ We propose a novel self-supervised learning method that explicitly captures the relationship between local features and whole structure in point clouds by combining physical laws with geometric modeling. This is also the first self-supervised learning method that leverages the mechanism of force propagation to model the relationship between local and whole structures of objects.
@@
$\bullet$ We demonstrate the effectiveness of our method through extensive evaluations on multiple downstream tasks, including object classification and segmentation. Additionally, the incorporation of physical information enhances the understanding of 3D point clouds.
II Related Work
II-A Point Cloud Representation Learning
In recent years, the advancement of deep learning techniques has enabled direct utilization of point cloud data for learning intrinsic feature representations of shapes. Early methods, such as PointNet [5], used shared MLPs and max pooling operations to learn global features of point clouds. Subsequently, PointNet++ [6] introduced a hierarchical structure and local feature learning mechanisms to address the issue of PointNet’s insensitivity to local features in point clouds. DGCNN [7] further improves local feature extraction by dynamically constructing graphs based on point neighborhoods. Similarly, recent studies [21, 22, 23] have focused on learning local feature information from point clouds, aiming to extract as many semantic features as possible. However, the methods often treat point clouds as static geometric distributions, overlooking the explicit relationship between local and whole structures. This limitation, especially in the case of object deformation, makes it difficult to capture the impact of local changes on the whole structure. Against this backdrop, implicit field modeling [24, 25] has been gradually introduced into point cloud learning, representing the geometric structure of point clouds through continuous functions. This representation inherently provides spatial continuity and resolution adaptivity, making it well-suited for capturing the continuous relationship between local and whole structures. It also offers a stronger geometric foundation for subsequent unsupervised modeling.
II-B Unsupervised Point Cloud Representation Learning
Nowadays, numerous unsupervised learning methods have emerged. Generative methods learn features through self-reconstruction [26, 8, 27, 28, 11, 29]. For example, HC-VAE [27] introduces a hierarchical consistency mechanism to progressively generate high-quality point clouds. FoldingNet [28] reconstructs arbitrary point clouds using a graph-based encoder and a folding-based decoder. Recently, recovering missing parts from masked inputs [9, 10, 8, 30] has also become a prominent research focus. MaskSurf [9] designs a transformer-based encoder-decoder network to estimate the position of the masking point and its corresponding normal vector. Point-DAE [8] reconstructs various corrupted point clouds by decomposing the reconstruction of a complete point cloud into detailed local patches and a coarse global shape, thereby alleviating the issue of positional leakage in the reconstruction process.
Discriminative methods guide the model to learn meaningful feature representations by designing tasks that focus on the feature differences between samples. Some methods employ pretext tasks to learn point cloud representations [31, 32, 14, 33, 34, 35]. Jigsaw3D [31] uses 3D puzzles as a self-supervised learning task, allowing the model to learn how to correctly assemble object parts. Poursaeed et al. [36] define the estimation of point cloud rotation angles as a pretext task. Recently, contrastive learning methods [15, 37, 14] have been widely used in 3D self-supervised learning, achieving state-of-the-art performance. PointContrast [14] performs point-level invariant mapping on two different directional views of the same point cloud. It is the first unified framework to study the contrastive paradigm for 3D representation learning. STRL [15] is an extension of the method [38] on 3D point clouds to learn rich representational information by updating between online and target networks. PointSmile [17] guides network learning by maximizing the mutual information between two point clouds. In addition, some researchers have focused on incorporating cross-model [39, 40, 41, 16, 42, 43, 44] methods into discriminative methods. However, these methods often overlook the influence of local structures on whole shape responses and typically rely on large numbers of positive-negative sample pairs and complex strategies, resulting in a learning process that lacks physical interpretability.
@@ II-B Unsupervised Point Cloud Representation Learning
In this work, we incorporate a physics-driven mechanism into the data-driven method that explicitly models deformation propagation from local forces to whole structures, enabling physically consistent and interpretable point cloud understanding. This also leads to more robust and generalizable features for downstream tasks.
III Method
@@
The framework of our method (Figure 2) consists of two main modules. The first module is the implicit feature learning module, where the point cloud is input into an autoencoder to learn its implicit features. The implicit field represents the point cloud as a continuous surface, and by leveraging the distance function from a query point to the nearest surface, it can uniformly handle objects of arbitrary shapes, helping the network better learn the whole shape features of the object. The second module is the physical information awareness module, which captures local deformation behaviors under external forces by constructing tetrahedral meshes and introducing deformation mechanisms from elasticity theory. This module not only learns the physical characteristics of local responses, but also embeds local deformation features into a unified geometric representation space via a shared encoder, thereby enabling mutual enhancement between whole shape representation and local physical perception. The proposed dual-task self-supervised framework combines the interpretability of physics-based modeling with the strong representational capacity of data-driven methods, enabling more effective learning of the geometric representations of real-world objects.
III-A Implicit Feature Learning
@@
For a point cloud dataset $\mathcal{D}=\left\{{{P}_{i}}\right\}_{i=1}^{\left|\mathcal{D}\right|}$, we use encoder ${\Phi}_{\theta}$} and decoder ${\Psi}_{\eta}$} to learn the implicit features of the point cloud ${P_{i}}=\left\{{p_{i}^{j}\in{\mathbb{R}^{3}}|j=1,2,...,n}\right\}$. The encoder ${\Phi}_{\theta}$} takes the $P_{i}$} as input and maps it to an $m$-dimensional latent vecoter $z_{\eta}$}, represented as ${{\Phi}_{\theta}}:{{\mathbb{R}}^{n\times 3}}\to{{\mathbb{R}}^{m}},m\ll n$, where we use the network backbones of PointNet [5] or DGCNN [7] as ${\Phi}_{\theta}$} to learn the point cloud features, respectively. Subsequently, we concatenate latent vector $z_{\eta}$} with an arbitrary query point $q$ in 3D space to obtain feature vector $z_{\eta}^{q}={z_{\eta}}\oplus q$. Then, the decoder ${\Psi}_{\eta}$} takes $z_{\eta}^{q}$ as input and output an implicit function $\mathbb{S}={{\Psi}_{\eta}}\left(q,{{\Phi}_{\theta}}\left({{P}_{i}}\right)\right)=\left\{{\delta_{i}}|{\delta_{i}}\left({z_{\eta}^{q}}\right)\in\mathbb{R},\forall q\in{{\mathbb{R}}^{3}}\right\}$. ${\delta_{i}}\left({z_{\eta}^{q}}\right)$ represents the shortest distance from the query point $q$ to the point cloud $P_{i}$}. The decoder ${\Psi}_{\eta}$} consists of 3-layer of MLP and outputs a 1-dimensional feature vector.
@@
${\theta^{*}},{\eta^{*}}=\mathop{\arg\min}\limits_{\theta,\eta}\left({\delta_{i}(z_{\eta}^{q})),{\delta_{gt}}\left(q\right)}\right),$ (1)
@@
${{\cal L}_{imp}}=\frac{1}{{|\mathcal{D}|K}}\sum\limits_{i=1}^{|\mathcal{D}|}{\sum\limits_{q\in Q}^{K}{\left|{\left|{{\delta_{i}}(z_{\eta}^{q})}\right|-{\delta_{gt}(q)}}\right|}},$ (2)
@@
III-B Physical Information Awareness
For a point cloud ${P}_{i}$}, we first use 3D Delaunay [46] to convert the ${P}_{i}$} into an object composed of multiple tetrahedra. It is an extension of Delaunay triangulation (which generates triangular meshes in 2D) to 3D space, producing a mesh composed of tetrahedral elements. Compared to hexahedral meshes, tetrahedral meshes better fit free-form objects and support complex topological structures, providing the volumetric elements required for physical modeling, such as deformation gradients and stress tensor computation. Meanwhile, directly using point coordinates lacks structural information, making it difficult to establish the strain–stress relationships necessary for constructing the subsequent physics-informed loss ${\mathcal{L}}_{pi}$. The advantages of using tetrahedral structures over point-based representations will be further validated in the ablation studies. The meshing process of the point cloud is illustrated in Figure 3.
Subsequently, the tetrahedral mesh data is processed by the mesh processor $\mathrm{T}$, which encodes the local geometric and topological structure of each tetrahedron into a unified format suitable for input to the encoder ${\Phi}_{\theta}$}, where the mesh processor $\mathrm{T}$ consists of 3-layer of convolutional networks, serving as a bridge between the mesh data and encoder ${\Phi}_{\theta}$}. The encoder ${\Phi}_{\theta}$} maps the processed data to the $m$-dimensional latent space to obtain the latent vector $z_{phys}$}. $z_{phys}$} is concatenated with the deformation parameters $d$ as the input $z_{phy}^{d}$ to decoder ${\Psi}_{phys}$}, enabling the model to learn local-to-whole deformation patterns under varying physical conditions during training. The $d$ includes standard physical parameters commonly used in static linear elasticity theory, such as Lamé constants $\lambda$ and $\mu$, elastic modulus $E$, Poisson’s ratio $\nu$, and external forces $f$. These parameters have clear physical meanings and can effectively characterize stress responses of typical elastic materials under small deformations. Considering the differences in point cloud density and actual deformation scale, we further apply an adaptive adjustment to the parameters in $d$ based on the size of the generated tetrahedral mesh and the initial deformation magnitude. This reduces reliance on manual configuration and enhances the robustness and generalization ability of the proposed method. Decoder ${\Psi}_{phys}$} is composed of 5-layer of MLP, outputting a feature vector of size $n\times 3$, where $n$ is the number of input points in the point cloud.